# The SPHEREx Satellite Mission

### v1

@@ 3.3 Spatial Resolution
Using the simulation tools described in Stickley et al. (2016), we studied photometry on point sources in the presence of surrounding stars and galaxies below the SPHEREx detection level with variously sized pixels and PSFs. These calculations confirmed that photometry is limited by instrumental (photon noise dominated) sensitivity for the all-sky survey, and the effects of confusion are small. Consequently, reducing either the PSF or the pixel size has little impact on the number of measured redshifts. Indeed, we observed that a smaller PSF with a fixed 62 pixel size improved photometric sensitivity, though very modestly, by slightly reducing the effective number of pixels observing a point source.
Spectral confusion adds wavelength-coherent variations that depend on the pixel scale and the depth of sources extracted for photometry (Huai et al., 2025). For the all-sky survey, 62 pixels limit the density of targeted sources due to beam overlap increasing noise, but spectral confusion itself is small compared with instrument noise. At the higher sensitivity in the deep fields, spectral confusion may exceed instrument noise. For the galaxy formation science theme in §2.2, the requirement is to mask or otherwise remove point sources from the map. For this science, 62 pixels suffice, because improved spatial resolution only helps remove the faintest sources. Finally, Huai et al. (2025) quantify the photometric depth and galaxy redshifts that can be reliably extracted from the deep field due to the effect of spectral confusion.
@@ 3.4.1 Gain Stability
We intend to correct gain variations to % over 2 years for inflationary cosmology science. The instrument controls gain stability by addressing the detectors with stabilized biases and references that are on thermally regulated oven-controlled regions in the warm electronics circuit boards, and by thermally regulating the H2RG detectors. The WISE mission shows remarkable stability with its H1RG arrays (Cutri et al., 2018), with variation in amplitude over 8 years after correcting for drifts in the uncontrolled detector temperatures. SPHEREx will monitor gain stability by measuring the relative flux of 100,000 bright stars in the deep fields. The deep fields are generally observed every orbit, and the stars provide ample statistical sensitivity to track relative gain changes in every readout channel.
@@ 3.4.2 Noise Bias
If uncorrected, variations in sensitivity on the sky lead to spurious clustering due to redshift non-uniformity. Because SPHEREx is photon noise limited, sensitivity variations are driven by spatial variations in sky brightness, primarily due to variations in Zodiacal light. We plan to mitigate the amplitude of the raw effect on spurious clustering (Doré et al., 2014) by a factor of 20 by injecting artificial sources into the data pipeline (Huff et al., 2014; Suchyta et al., 2016; Everett et al., 2022). Since the artificial sources are observed with noise realizations of real data, they accurately track the variations in redshift number density due to variations in sensitivity. We also model the detector noise with photocurrent (Nguyen et al., 2025), and record daily image pairs taken of the same field to provide an accurate in-flight assessment of the noise behavior per pixel.
@@ 3.4.4 Stray Light
SPHEREx uses an off-axis telescope design to control stray light on scales ranging from the 6″ PSF core to tens of degrees. An extended PSF produces ‘halos’ that are problematic around bright sources. The design reduces the extended PSF by controlling mirror roughness, particulate contamination, and multiple reflections between the LVFs and detectors. The extension is well measured in flight data and can be mitigated by a combination of masking and numerical subtraction. The response to bright stars outside the field of view is controlled by a series of baffles around the mirrors and detectors. Finally, far-angle response to the Earth and Galactic plane are controlled by the telescope design. We have carried out a series of in-flight measurements to assess the level of systematic error that results in the deep field maps used in the galaxy formation science theme. We will further check susceptibility by comparing subsets of the deep maps, made under different stray light conditions, for statistical consistency. A detailed analysis of the design and the in-flight measurements will be presented in a future publication (Dowell et al., 2026).
@@ 3.4.5 Dark Current
Detector dark current produces a systematic error for galaxy formation science by leaving a residual pattern in the deep field maps. The dark current levels reported in Table 3 were measured in the laboratory under flight-like conditions, using the flight cadence and regular resets in 32-channel output mode. We simulated deep field mosaics from these lab dark current images, and find a residual signal in a synthesized broad 20 % band at 1.2 m at the level of at an angular scale of . To reach our target level for this systematic, we plan to measure and subtract dark current to a modest accuracy of % using the offsets produced by the Zodiacal flat-field estimator described in §3.4.3. We have already produced dark current estimates using this method on early flight data and find general agreement with the laboratory measurements. Once the deep fields are fully observed, we will construct null tests with chosen combinations of images (Zemcov et al., 2014) to quantify the level of systematic error that arises from dark current and other mechanisms.
@@ 4.2 Survey Design
The SPHEREx survey (Bryan et al., 2025) is designed to cover the entire sky, observing every point on the sky in each of its 102 spectral channels, at least twice per year. We define spectral channels in intervals given by , where is the LVF transmission normalized to unity at peak. The required spectral range and resolving power set the minimum number of spectral channels. However, to simplify the LVF fabrication we fix the resolving power per band, and to simplify the survey design we require that all the spectral channels subtend the same angular size on the focal plane. Our design thus has 17 spectral channels (see §4.3.3 and Figure 12) for each of the 6 bands in Table 2. In the survey design, we consider a spectral channel to be measured when the observed wavelength falls within the geometrically-defined region. Note that each object will actually be observed at a unique set of wavelengths, and will typically have more than 102 wavelength observations in each survey. Because the interval of a spectral channel does not Nyquist sample the spectral response of the LVF, we shift the targets by half a spectral channel, i.e. , between the first and third, and second and fourth surveys.
@@ 4.3 Instrument Design
As a NASA Explorer satellite, SPHEREx is designed to carry out its scientific objectives with an economical implementation. We developed the observatory to carry out its survey using a spectrometer without moving parts, based on an entirely passive cooling system and a single-string architecture throughout the instrument and spacecraft. Further discussion of the instrument design, testing, and performance can be found in Korngut et al. (2025).
@@ 4.3.1 Cooling System
SPHEREx uses a multi-stage radiative cooling system to achieve its desired operating temperatures for the optical system and FPAs. The basic approach follows the thermal design of the Planck satellite, which used a 3-stage V-groove radiator to cool its radiator panels to 140, 90 and 46 K, with the telescope reaching 36 K (Planck Collaboration et al., 2011). However, Planck operated in a halo orbit about the second Earth-Sun Lagrange point (L2), which avoided emission from the Earth and naturally shaded the V-grooves from direct sunlight by the spacecraft body and solar panel.
Following Planck, SPHEREx has 3 V-groove panels located between the spacecraft and the telescope (see Figure 6). Being in low-Earth orbit, the design incorporates 3 conical “photon shields”, which serve as extensions of the 3 V-groove coolers. The photon shield cone angles are progressively increased from inner to outer to allow thermal photons to escape to space in a few specular reflections. The nested cones also protect the cold instrument from being illuminated by the Sun or the Earth. During observations, the Earth illuminates the inner photon shield, but at an angle where the light then reflects to cold space without intercepting the instrument. The telescope and focal planes are supported on low-conductivity bipods, which are thermally intercepted at each of the 3 V-grooves to reduce conducted power to the instrument. The thermally isolated telescope also radiates to space, effectively comprising a fourth radiation stage. Finally, the MWIR FPA requires the lowest temperature. It is cooled by a dedicated radiator panel, the fifth radiation stage, that is isolated from the telescope by bipods. Moore et al. (2025) will present further information on the design and in-flight performance of the SPHEREx thermal system (also see §5.2).
@@ 4.3.2 Optics
BAE assembled and aligned the telescope, and cryogenically tested the image quality before and after environmental vibration testing (Frater et al., 2023). The mirrors and housing were all fabricated from a single billet of Aluminum 6061 alloy in order to minimize misalignment at cryogenic temperature. The mirrors were fabricated with light-weighted back surfaces and coated with FSG98-protected gold to reduce reflection loss. After assembly, the measured spot sizes varied from 3.9 to 10.8 (SWIR) and 7 to 14.7 (MWIR) at ambient temperature, and from 4.4 to 17.4 (SWIR) at 58 K (the cryogenic performance was only measured in reflection off the DBS with optical interferometry). We attribute the increase, most evident at the extreme corners of bands 3 (see Figure 17), to alignment errors and mounting stress in the mirrors. The image quality, assuming ideal placement of the focal plane surface, is better than FWHM over most of the SWIR focal plane. In the extreme corners in band 3, the PSF becomes as large as 6″. The PSF in the MWIR focal plane increases up to due to diffraction at longer wavelengths.
@@ 4.3.3 Spectroscopy
SPHEREx carries out spectroscopy by placing high-efficiency LVFs over each of the 6 H2RG focal plane detector arrays. While this simple approach is unique to astrophysics, it was used in two NASA planetary missions (Reuter et al., 2008; Reuter & Simon-Miller, 2012) to rapidly map the surfaces of solar system objects in low-resolution infrared spectroscopy.
@@
The spectral response of the instrument was measured at operating temperature in the laboratory using a specialized cryostat equipped with a cooled integrating sphere and Winston horn arrangement to illuminate the focal surface uniformly. The integrating sphere was coupled to a laboratory grating spectrometer to measure the spectral response of every pixel in the focal plane, providing measurements of the band width, band center, and out-of-band blocking. We show the central wavelength measured over the 6 focal plane arrays in Figure 12. A full description of the measurement apparatus and results will be presented in Hui et al. (2025).
@@ 4.3.4 Focal Plane Assemblies
SPHEREx uses 6 Hawaii-2RG detector arrays (Blank et al., 2012), part of a family of detectors with a long heritage in space-based applications (Beletic, 2024), housed in the two FPAs. The SWIR FPA views the DBS in reflection, while the MWIR FPA views the DBS in transmission. The SWIR detectors were fabricated with a cut-off wavelength of 2.65 m while the MWIR detectors have a cut-off wavelength of 5.4 m. The properties of the detectors are summarized in Table 3. Note that each H2RG detector provides a 4 pixel-wide frame of non-photosensitive reference pixels that can be used to assess dark current and noise properties.
@@ 4.3.5 Readout Electronics
SPHEREx warm electronics (see Heaton et al. (2023) for a detailed description) control and read the detector arrays, digitize the data, perform sample-up-the-ramp (SUR; Garnett & Forrest (1993)) fitting to extract the detector current, compress the resulting data, and pass the output to the spacecraft computer for telemetry. The electronics also reads out instrument thermometry and controls the SWIR and MWIR temperatures. The readout electronics employ 6 identical readout boards, which provide bias and timing signals for each of the 6 detector arrays. The readout boards are operated by a central electronics board that communicates with the spacecraft computer. The bias voltages for the H2RGs are generated in an oven-controlled region of the readout board to minimize thermally-induced drifts.
@@
SPHEREx uses the SUR algorithm to reduce the data volume to fit within the available telemetry bandwidth. SUR provides a best-fit slope to measure the photocurrent using a constant sampling cadence every 1.5 s. The regular sampling provides additional flexibility, enabling the detection of transients and saturation on board. Our SUR algorithm (Zemcov et al., 2016) identifies saturation on bright sources, and for sources that do not saturate in a few reads, returns a best-fit slope over a restricted time period to give a larger dynamic range. The algorithm also searches for voltage jumps caused by energetic particles, to a chosen threshold, and returns a best-fit slope from data prior to the event. The algorithm returns flags identifying these special cases, and also a flag indicating if the flagged voltage jump occurred in the first half or second half of the exposure.
@@
Row-chopping is very effective in removing the effect of noise within a readout channel. However, the common-mode leaves a small drift that translates to a different offset current in each readout channel. We can average the 8 rows of photo-insensitive reference pixels available in the H2RG detector to measure and subtract the offsets in each channel. Laboratory tests in a dark test cryostat (Nguyen et al., 2025) using this method showed the low spatial frequency noise was below the expected level of photon noise in bands 1 - 4 for spatial wavenumber (or multipole for our 62 pixel scale). Note the photon noise level in Figure 9 of Nguyen et al. (2025) is from an early estimate of optical efficiency, and should be increased by a factor of 4 (in units) to match the in-flight photocurrent. The tests also showed that subtracting the median of the optical pixels within a readout channel further improves the performance, as the channel offsets can be measured more accurately (even with photon noise) than by averaging the comparatively small number of reference pixels.
@@ 4.5 Mission Operations
Mission operations are conducted at JPL from the Earth Orbiter Mission Operations Center using ground systems adapted from the WISE/NEOWISE missions. The Survey Planning Software (Bryan et al., 2025), developed by Arizona State University, is run for the next planning period using the latest orbit prediction to provide an accurate forecast of Earth and Moon avoidance constraints. The team generates command sequences twice a week from the survey plan, and checks constraints independently. During a typical day, the spacecraft transmits science data through several 5 minute passes. The spacecraft points during these maneuvers to maintain the telemetry link while adhering to the avoidance constraints. No imaging data is taken during these downlinks. The team issues commands and receives telemetry via the S-band link. The mission operations team typically downlinks 20 GB of science data each day at 600 Mbps via the Ka-band link. All communications are routed through NASA’s NSN ground stations, including those at Fairbanks in Alaska, Punta Arenas in Chile, Svalbard in Norway and the Troll Satellite Station in Antarctica.

### v2

@@ 3.3 Spatial Resolution
Using the simulation tools described in Stickley16, we studied photometry on point sources in the presence of surrounding stars and galaxies below the SPHEREx detection level with variously sized pixels and PSFs. These calculations confirmed that photometry is limited by instrumental (photon noise dominated) sensitivity for the all-sky survey, and the effects of confusion are small. Consequently, reducing either the PSF or the pixel size has little impact on the number of measured redshifts. Indeed, we observed that a smaller PSF with a fixed 62 pixel size improved photometric sensitivity, though very modestly, by slightly reducing the effective number of pixels observing a point source.
Spectral confusion adds wavelength-coherent variations that depend on the pixel scale and the depth of sources extracted for photometry (Huai25). For the all-sky survey, 62 pixels limit the density of targeted sources due to beam overlap increasing noise, but spectral confusion itself is small compared with instrument noise. At the higher sensitivity in the deep fields, spectral confusion may exceed instrument noise. For the galaxy formation science theme in §2.2, the requirement is to mask or otherwise remove point sources from the map. For this science, 62 pixels suffice, because improved spatial resolution only helps remove the faintest sources. Finally, Huai25 quantify the photometric depth and galaxy redshifts that can be reliably extracted from the deep field due to the effect of spectral confusion.
@@ 3.4.1 Gain Stability
We intend to correct gain variations to % over 2 years for inflationary cosmology science. The instrument controls gain stability by addressing the detectors with stabilized biases and references that are on thermally regulated oven-controlled regions in the warm electronics circuit boards, and by thermally regulating the H2RG detectors. The WISE mission shows remarkable stability with its H1RG arrays (cutri18), with variation in amplitude over 8 years after correcting for drifts in the uncontrolled detector temperatures. SPHEREx will monitor gain stability by measuring the relative flux of 100,000 bright stars in the deep fields. The deep fields are generally observed every orbit, and the stars provide ample statistical sensitivity to track relative gain changes in every readout channel.
@@ 3.4.2 Noise Bias
If uncorrected, variations in sensitivity on the sky lead to spurious clustering due to redshift non-uniformity. Because SPHEREx is photon noise limited, sensitivity variations are driven by spatial variations in sky brightness, primarily due to variations in Zodiacal light. We plan to mitigate the amplitude of the raw effect on spurious clustering (Dore14) by a factor of 20 by injecting artificial sources into the data pipeline (huff14; suchyta16; Everett:2020). Since the artificial sources are observed with noise realizations of real data, they accurately track the variations in redshift number density due to variations in sensitivity. We also model the detector noise with photocurrent (Nguyen25), and record daily image pairs taken of the same field to provide an accurate in-flight assessment of the noise behavior per pixel.
@@ 3.4.4 Stray Light
SPHEREx uses an off-axis telescope design to control stray light on scales ranging from the 6″ PSF core to tens of degrees. An extended PSF produces ‘halos’ that are problematic around bright sources. The design reduces the extended PSF by controlling mirror roughness, particulate contamination, and multiple reflections between the LVFs and detectors. The extension is well measured in flight data and can be mitigated by a combination of masking and numerical subtraction. The response to bright stars outside the field of view is controlled by a series of baffles around the mirrors and detectors. Finally, far-angle response to the Earth and Galactic plane are controlled by the telescope design. We have carried out a series of in-flight measurements to assess the level of systematic error that results in the deep field maps used in the galaxy formation science theme. We will further check susceptibility by comparing subsets of the deep maps, made under different stray light conditions, for statistical consistency. A detailed analysis of the design and the in-flight measurements will be presented in a future publication (Dowell25).
@@ 3.4.5 Dark Current
Detector dark current produces a systematic error for galaxy formation science by leaving a residual pattern in the deep field maps. The dark current levels reported in Table 3 were measured in the laboratory under flight-like conditions, using the flight cadence and regular resets in 32-channel output mode. We simulated deep field mosaics from these lab dark current images, and find a residual signal in a synthesized broad 20 % band at 1.2 m at the level of at an angular scale of . To reach our target level for this systematic, we plan to measure and subtract dark current to a modest accuracy of % using the offsets produced by the Zodiacal flat-field estimator described in §3.4.3. We have already produced dark current estimates using this method on early flight data and find general agreement with the laboratory measurements. Once the deep fields are fully observed, we will construct null tests with chosen combinations of images (zemcov14) to quantify the level of systematic error that arises from dark current and other mechanisms.
@@ 4.2 Survey Design
The SPHEREx survey (bryan25) is designed to cover the entire sky, observing every point on the sky in each of its 102 spectral channels, at least twice per year. We define spectral channels in intervals given by , where is the LVF transmission normalized to unity at peak. The required spectral range and resolving power set the minimum number of spectral channels. However, to simplify the LVF fabrication we fix the resolving power per band, and to simplify the survey design we require that all the spectral channels subtend the same angular size on the focal plane. Our design thus has 17 spectral channels (see §4.3.3 and Figure 12) for each of the 6 bands in Table 2. In the survey design, we consider a spectral channel to be measured when the observed wavelength falls within the geometrically-defined region. Note that each object will actually be observed at a unique set of wavelengths, and will typically have more than 102 wavelength observations in each survey. Because the interval of a spectral channel does not Nyquist sample the spectral response of the LVF, we shift the targets by half a spectral channel, i.e. , between the first and third, and second and fourth surveys.
@@ 4.3 Instrument Design
As a NASA Explorer satellite, SPHEREx is designed to carry out its scientific objectives with an economical implementation. We developed the observatory to carry out its survey using a spectrometer without moving parts, based on an entirely passive cooling system and a single-string architecture throughout the instrument and spacecraft. Further discussion of the instrument design, testing, and performance can be found in Korngut25.
@@ 4.3.1 Cooling System
SPHEREx uses a multi-stage radiative cooling system to achieve its desired operating temperatures for the optical system and FPAs. The basic approach follows the thermal design of the Planck satellite, which used a 3-stage V-groove radiator to cool its radiator panels to 140, 90 and 46 K, with the telescope reaching 36 K (Planck_thermal11). However, Planck operated in a halo orbit about the second Earth-Sun Lagrange point (L2), which avoided emission from the Earth and naturally shaded the V-grooves from direct sunlight by the spacecraft body and solar panel.
Following Planck, SPHEREx has 3 V-groove panels located between the spacecraft and the telescope (see Figure 1). Being in low-Earth orbit, the design incorporates 3 conical “photon shields”, which serve as extensions of the 3 V-groove coolers. The photon shield cone angles are progressively increased from inner to outer to allow thermal photons to escape to space in a few specular reflections. The nested cones also protect the cold instrument from being illuminated by the Sun or the Earth. During observations, the Earth illuminates the inner photon shield, but at an angle where the light then reflects to cold space without intercepting the instrument. The telescope and focal planes are supported on low-conductivity bipods, which are thermally intercepted at each of the 3 V-grooves to reduce conducted power to the instrument. The thermally isolated telescope also radiates to space, effectively comprising a fourth radiation stage. Finally, the MWIR FPA requires the lowest temperature. It is cooled by a dedicated radiator panel, the fifth radiation stage, that is isolated from the telescope by bipods. Moore25 will present further information on the design and in-flight performance of the SPHEREx thermal system (also see §5.2).
@@ 4.3.2 Optics
BAE assembled and aligned the telescope, and cryogenically tested the image quality before and after environmental vibration testing (Frater23). The mirrors and housing were all fabricated from a single billet of Aluminum 6061 alloy in order to minimize misalignment at cryogenic temperature. The mirrors were fabricated with light-weighted back surfaces and coated with FSG98-protected gold to reduce reflection loss. After assembly, the measured spot sizes varied from 3.9 to 10.8 (SWIR) and 7 to 14.7 (MWIR) at ambient temperature, and from 4.4 to 17.4 (SWIR) at 58 K (the cryogenic performance was only measured in reflection off the DBS with optical interferometry). We attribute the increase, most evident at the extreme corners of bands 3 (see Figure 17), to alignment errors and mounting stress in the mirrors. The image quality, assuming ideal placement of the focal plane surface, is better than FWHM over most of the SWIR focal plane. In the extreme corners in band 3, the PSF becomes as large as 6″. The PSF in the MWIR focal plane increases up to due to diffraction at longer wavelengths.
@@ 4.3.3 Spectroscopy
SPHEREx carries out spectroscopy by placing high-efficiency LVFs over each of the 6 H2RG focal plane detector arrays. While this simple approach is unique to astrophysics, it was used in two NASA planetary missions (Reuter08; Reuter12) to rapidly map the surfaces of solar system objects in low-resolution infrared spectroscopy.
@@
The spectral response of the instrument was measured at operating temperature in the laboratory using a specialized cryostat equipped with a cooled integrating sphere and Winston horn arrangement to illuminate the focal surface uniformly. The integrating sphere was coupled to a laboratory grating spectrometer to measure the spectral response of every pixel in the focal plane, providing measurements of the band width, band center, and out-of-band blocking. We show the central wavelength measured over the 6 focal plane arrays in Figure 12. A full description of the measurement apparatus and results will be presented in Hui25.
@@ 4.3.4 Focal Plane Assemblies
SPHEREx uses 6 Hawaii-2RG detector arrays (Blank12), part of a family of detectors with a long heritage in space-based applications (Beletic24), housed in the two FPAs. The SWIR FPA views the DBS in reflection, while the MWIR FPA views the DBS in transmission. The SWIR detectors were fabricated with a cut-off wavelength of 2.65 m while the MWIR detectors have a cut-off wavelength of 5.4 m. The properties of the detectors are summarized in Table 3. Note that each H2RG detector provides a 4 pixel-wide frame of non-photosensitive reference pixels that can be used to assess dark current and noise properties.
@@ 4.3.5 Readout Electronics
SPHEREx warm electronics (see Heaton23 for a detailed description) control and read the detector arrays, digitize the data, perform sample-up-the-ramp (SUR; Garnett93) fitting to extract the detector current, compress the resulting data, and pass the output to the spacecraft computer for telemetry. The electronics also reads out instrument thermometry and controls the SWIR and MWIR temperatures. The readout electronics employ 6 identical readout boards, which provide bias and timing signals for each of the 6 detector arrays. The readout boards are operated by a central electronics board that communicates with the spacecraft computer. The bias voltages for the H2RGs are generated in an oven-controlled region of the readout board to minimize thermally-induced drifts.
@@
SPHEREx uses the SUR algorithm to reduce the data volume to fit within the available telemetry bandwidth. SUR provides a best-fit slope to measure the photocurrent using a constant sampling cadence every 1.5 s. The regular sampling provides additional flexibility, enabling the detection of transients and saturation on board. Our SUR algorithm (Zemcov16) identifies saturation on bright sources, and for sources that do not saturate in a few reads, returns a best-fit slope over a restricted time period to give a larger dynamic range. The algorithm also searches for voltage jumps caused by energetic particles, to a chosen threshold, and returns a best-fit slope from data prior to the event. The algorithm returns flags identifying these special cases, and also a flag indicating if the flagged voltage jump occurred in the first half or second half of the exposure.
@@
Row-chopping is very effective in removing the effect of noise within a readout channel. However, the common-mode leaves a small drift that translates to a different offset current in each readout channel. We can average the 8 rows of photo-insensitive reference pixels available in the H2RG detector to measure and subtract the offsets in each channel. Laboratory tests in a dark test cryostat (Nguyen25) using this method showed the low spatial frequency noise was below the expected level of photon noise in bands 1 - 4 for spatial wavenumber (or multipole for our 62 pixel scale). Note the photon noise level in Figure 9 of Nguyen25 is from an early estimate of optical efficiency, and should be increased by a factor of 4 (in units) to match the in-flight photocurrent. The tests also showed that subtracting the median of the optical pixels within a readout channel further improves the performance, as the channel offsets can be measured more accurately (even with photon noise) than by averaging the comparatively small number of reference pixels.
@@ 4.5 Mission Operations
Mission operations are conducted at JPL from the Earth Orbiter Mission Operations Center using ground systems adapted from the WISE/NEOWISE missions. The Survey Planning Software (bryan25), developed by Arizona State University, is run for the next planning period using the latest orbit prediction to provide an accurate forecast of Earth and Moon avoidance constraints. The team generates command sequences twice a week from the survey plan, and checks constraints independently. During a typical day, the spacecraft transmits science data through several 5 minute passes. The spacecraft points during these maneuvers to maintain the telemetry link while adhering to the avoidance constraints. No imaging data is taken during these downlinks. The team issues commands and receives telemetry via the S-band link. The mission operations team typically downlinks 20 GB of science data each day at 600 Mbps via the Ka-band link. All communications are routed through NASA’s NSN ground stations, including those at Fairbanks in Alaska, Punta Arenas in Chile, Svalbard in Norway and the Troll Satellite Station in Antarctica.